OpenStudy (anonymous):

write the equation of a line parallel to the given line but passing through the given point y=-3x-6;( -1,5)

OpenStudy (anonymous):

I do not understand where and how you got that answer

OpenStudy (anonymous):

For a line to be parallel, it means that it has the same gradient as the given line. In this case the gradient m=-3. Then, you know that the equation of a line is: y−y0=m∗(x−x0) in this case, y0=5; x0=-1. then the equation simplifies to y=-3x+2, after substitution. (y-5=-3*(x-(-1)) => y=-3*(x+1)+5 => y=-3x-3+5 => y=-3x+2)

OpenStudy (anonymous):

ok so what is the answer

OpenStudy (anonymous):

and thatnk you very much for helping me

OpenStudy (anonymous):

y=-3x+2

OpenStudy (anonymous):

No bother : )

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

yes thank you again